हिंदी

Solve the following: A and B throw a die alternatively till one of them gets a 3 and wins the game. Find the respective probabilities of winning. (Assuming A begins the game)

Advertisements
Advertisements

प्रश्न

Solve the following:

A and B throw a die alternatively till one of them gets a 3 and wins the game. Find the respective probabilities of winning. (Assuming A begins the game)

योग
Advertisements

उत्तर

Let A ≡ the event that player A gets 3

B ≡ the event that player B gets 3

When a die is thrown, the probability of getting 3 is `1/6`

∴ P(A) = `1/6`, P(A') = 1 – P(A) = `5/6`

Similarly, P(B) = `1/6`, P(B') = `5/6`

It is given that A starts the game and he will win in the following mutually exclusive ways.

(i) A happens i.e., A wins at the first throw.

(ii) A' n B' n A happens i.e., A wins in the third throw when A and B fail in the first and second throws.

(iii) A' n B' n A' n B' n A happens i.e., A wins at the fifth throw when A and B fail at the 1st, 2nd, 3rd, 4th throw, and so on.

∴ P(A wins)= P(A) + P(A' ∩ B' ∩ A) + P(A' ∩ B' ∩ A' ∩ B' ∩ A) + .......

= P(A) + P(A') · P(B') · P(A) + P(A') · P(B') · P(A') · P(B') · P(A) + .......

= `1/6 + 5/6 xx 5/6 + 1/6 + 5/6 xx 5/6 xx 5/6 xx 5/6 xx 1/6` + ...

= `1/6[1 + (5/6)^2 + (5/6)^4 + ...]`

= `1/6[1/(1 - (5/6)^2)]`  ...[it is an infinite geometric series with a = 1, r = `25/36`]

= `1/6[1/((11/36))]`

= `6/11`

P(B wins)= 1 – P(A wins)

= `1 - 6/11 = 5/11`

shaalaa.com
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 9: Probability - Miscellaneous Exercise 9 [पृष्ठ २१४]

APPEARS IN

बालभारती Mathematics and Statistics (Arts and Science) Part 1 [English] Standard 11 Maharashtra State Board
अध्याय 9 Probability
Miscellaneous Exercise 9 | Q II. (18) | पृष्ठ २१४

वीडियो ट्यूटोरियलVIEW ALL [2]

संबंधित प्रश्न

A bag contains 4 balls. Two balls are drawn at random (without replacement) and are found to be white. What is the probability that all balls in the bag are white?


If `P(A)  = 3/5 and P(B) = 1/5` , find P (A ∩ B) if A and B are independent events.


One card is drawn at random from a well-shuffled deck of 52 cards. In which of the following case is the events E and F independent?

E : ‘the card drawn is a king or queen’

F : ‘the card drawn is a queen or jack’


A speaks the truth in 60% of the cases, while B is 40% of the cases. In what percent of cases are they likely to contradict each other in stating the same fact?


A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, exactly two students solve the problem?


A, B, and C try to hit a target simultaneously but independently. Their respective probabilities of hitting the target are `3/4, 1/2` and `5/8`. Find the probability that the target

  1. is hit exactly by one of them
  2. is not hit by any one of them
  3. is hit
  4. is exactly hit by two of them

The probability that a man who is 45 years old will be alive till he becomes 70 is `5/12`. The probability that his wife who is 40 years old will be alive till she becomes 65 is `3/8`. What is the probability that, 25 years hence,

  1. the couple will be alive
  2. exactly one of them will be alive
  3. none of them will be alive
  4. at least one of them will be alive

A bag contains 3 yellow and 5 brown balls. Another bag contains 4 yellow and 6 brown balls. If one ball is drawn from each bag, what is the probability that, both the balls are of the same color?


A bag contains 3 yellow and 5 brown balls. Another bag contains 4 yellow and 6 brown balls. If one ball is drawn from each bag, what is the probability that, the balls are of different color?


Solve the following:

If P(A) = `"P"("A"/"B") = 1/5, "P"("B"/"A") = 1/3` the find `"P"("B'"/"A'")`


Solve the following:

Let A and B be independent events with P(A) = `1/4`, and P(A ∪ B) = 2P(B) – P(A). Find `"P"("A"/"B")`


Solve the following:

Let A and B be independent events with P(A) = `1/4`, and P(A ∪ B) = 2P(B) – P(A). Find `"P"("B'"/"A")`


Solve the following:

Find the probability that a year selected will have 53 Wednesdays


The probability of simultaneous occurrence of at least one of two events A and B is p. If the probability that exactly one of A, B occurs is q, then prove that P(A′) + P(B′) = 2 – 2p + q.


Let A and B be two independent events. Then P(A ∩ B) = P(A) + P(B)


Three events A, B and C are said to be independent if P(A ∩ B ∩ C) = P(A) P(B) P(C).


For a loaded die, the probabilities of outcomes are given as under:
P(1) = P(2) = 0.2, P(3) = P(5) = P(6) = 0.1 and P(4) = 0.3. The die is thrown two times. Let A and B be the events, ‘same number each time’, and ‘a total score is 10 or more’, respectively. Determine whether or not A and B are independent.


Refer to Question 1 above. If the die were fair, determine whether or not the events A and B are independent.


The probability that at least one of the two events A and B occurs is 0.6. If A and B occur simultaneously with probability 0.3, evaluate `"P"(bar"A") + "P"(bar"B")`


If A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A/B) = `1/4`, P(A' ∩ B') equals ______.


If A and B are two events and A ≠ Φ, B ≠ Φ, then ______.


If the events A and B are independent, then P(A ∩ B) is equal to ______.


Let P(A) > 0 and P(B) > 0. Then A and B can be both mutually exclusive and independent.


Two independent events are always mutually exclusive.


If A and B are two independent events then P(A and B) = P(A).P(B).


If A and B are independent, then P(exactly one of A, B occurs) = P(A)P(B') + P(B)P(A') 


One card is drawn at random from a well-shuffled deck of 52 cards. In which of the following case is the events E and F independent?

E : ‘the card drawn is a spade’

F : ‘the card drawn is an ace’


Let E1 and E2 be two independent events. Let P(E) denotes the probability of the occurrence of the event E. Further, let E'1 and E'2 denote the complements of E1 and E2, respectively. If P(E'1 ∩ E2) = `2/15` and P(E1 ∩ E'2) = `1/6`, then P(E1) is


If A, B are two events such that `1/8 ≤ P(A ∩ B) ≤ 3/8` then


Two events 'A' and 'B' are said to be independent if


Two cards are drawn at random and without replacement from a pack of 52 playing cards. Find the probability that both the cards are black.


The probability of the event A occurring is `1/3` and of the event B occurring is `1/2`. If A and B are independent events, then find the probability of neither A nor B occurring.


What are independent events in probability?


If E and F are independent, which pair is also independent?


For \[E=\{3,6\}\] and \[F=\{2,4,6\}\], what is \[E\cap F\]?


For \[E=\{3,6\}\] when a die is thrown once, what is \[P(E)\]?


For \[F=\{2,4,6\}\] when a die is thrown once, what is \[P(F)\]?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×